# Max Getter Chair for Geometry and Analysis

### v2

@@ Abstract
Index Terms:
Deep learning, representation learning, feature learning, scattering transform, information retention, energy propagation, neural networks.
@@ II-A A general model for feature extraction
[…]
$Sf:=\left\{\text{}S[p]f\leavevmode\nobreak\ :\leavevmode\nobreak\ \ell\geq 0,\leavevmode\nobreak\ p\in\mathcal{P}^{(\ell)}\text{}\right\},$
@@ II-B Energy conservation
The action of the $\ell$-th layer of a feedforward neural network [18] is typically described by $f^{(\ell)}=\sigma(W^{(\ell)}f^{(\ell-1)})$, where $W^{(\ell)}\in\mathbb{C}^{d^{(\ell)}\times d^{(\ell-1)}}$ are the (learned) weights including a potential bias term, $\sigma$ is a pointwise nonlinearity, and $f^{(\ell)}$ is the output of the $\ell$-th layer. Clearly, this fits into our framework, by setting $\mathcal{H}=\mathbb{C}^{d^{(0)}}$, $\mathcal{H}^{(\ell)}=\mathbb{C}^{d^{(\ell)}}$, ()=$\mathcal{L}^{(\ell)}=\left\{\text{}W^{(\ell)}\text{}\right\}$, $\sigma^{(\ell)}=\sigma$, for $\ell=1,\ldots,D$, as well as $A^{(\ell)}=0_{d^{(\ell)}\times d^{(\ell)}}$ for $\ell=0,\ldots,D-1$, and $A^{(D)}=\operatorname{id}_{\mathbb{C}^{d^{(D)}}}$. Assumption II.1 then boils down to the requirement that the singular values of $W^{(\ell)}$ are $\leq 1$. Linear feedforward neural networks [19] are a special instance thereof, taking pointwise identities for $\sigma$.
@@
Further, $S:\mathcal{H}\to\bigoplus_{\ell=0}^{\infty}\mathcal{H}^{(\ell)}$ is nonexpansive with $S(0)=0$, hence norm-decreasing.
@@
The infinite sum $\sum_{\ell=N}^{\infty}\sum_{p\in\mathcal{P}^{(\ell)}}\left\lVert\text{}S[p]f\text{}\right\rVert^{2}$ provides a natural measure to quantify the information contained in the layers of depth $\geq N$. This quantity is bounded above by $W_{N}(f)$. Assuming that $Sf$ is nontrivial, truncating the (possibly infinitely deep) feature extractor after the first $N$ layers looses at most $W_{N}(f)/\left\lVert\text{}Sf\text{}\right\rVert^{2}$ percent of the total energy that is distributed across the entire network, suggesting that a fast decay of this quantity is desirable (see also [7]). Rapid energy decay across layers is also provably closely related to the stability of scattering CNNs with respect to pertubations of the input [5, Corollary 2.15], [13, Theorem 5].
@@
$=\inf_{g\in\mathcal{R}^{(N)}:\left\lVert\text{}g\text{}\right\rVert=1}\left\lVert\text{}A^{(N)}g\text{}\right\rVert^{2},$ (4)
@@ II-C A generic rate for energy decay
where
$C^{(\ell)}:=\inf\limits_{g\in\mathcal{R}^{(\ell)}:\left\lVert\text{}g\text{}\right\rVert=1}\leavevmode\nobreak\ \sup\limits_{(\lambda,\eta)\in E^{(\ell)}}\leavevmode\nobreak\ \left|\left<\text{}\,g,\eta\,\text{}\right>\right|^{2}\cdot\lambda\geq 0.$ (6)
@@
While the case $E^{(\ell)}=\emptyset$ may occur in the very general case, we have $E^{(\ell)}\neq\emptyset$ in many practically relevant settings. In fact, the setting is inspired by that of scattering CNNs over general measure spaces [13], where $A^{(\ell)}=H^{t}$ is the diffusion operator defined in [13, Equ. (8),(10)] for some $t>0$, so $E^{(\ell)}=\{(g(\lambda_{k})^{2t},\varphi_{k})\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in I\}$ in the notation of [13].
@@
Assume that $C_{\mathcal{M}^{(\ell)}}:=\inf_{m}\leavevmode\nobreak\ \mu^{(\ell)}(m)>0$, where the infimum is taken over all sets $m\in\mathfrak{M}^{(\ell)}$ for which $\mu^{(\ell)}(m)>0$. Suppose further that $\sigma^{(\ell)}f\geq 0$ ($\mu^{(\ell)}$-a.e.) for all $f\in\mathcal{H}^{(\ell)}$ and that there exists a pair $(\lambda,\eta)\in E^{(\ell)}$ such that $\lambda^{1/2}\eta\geq C_{A^{(\ell)}}^{1/2}$ holds ($\mu^{(\ell)}$-a.e.) for a constant $C_{A^{(\ell)}}>0$. Then, we have
@@
In particular, if this holds for the first $\ell=1,\ldots,N$ layers of the neural network, then we have, for all $f\in\mathcal{H}$,
@@
Let us now continue the discussion from above regarding the concreteness and applicability of our results. First, note that the constraint on the measure space is satisfied for the counting measure, in which case $C_{\mathcal{M}^{(\ell)}}=1$. This may even be regarded as the most practically relevant measure space, having a feature engineer in mind who works with whatever feature extractor on a digital computer. Second, there is a whole class of feature extractors, for which there exists a constant eigenfunction, $\eta\equiv C_{A^{(\ell)}}^{1/2}=\mu^{(\ell)}(M^{(\ell)})^{-1/2}$. Indeed, for scattering CNNs defined via diffusion operators as in [13, Equ. (8)], the eigenfunction $\varphi_{0}$ of $\mathcal{L}$ (in the notation of [13]) is often constant [13, Remark 3], and we have $(1,\varphi_{0})\in E^{(\ell)}$ in the context of the above corollary; see also [13, Theorems 2, 5] and [5, Section 5.1], which are based on similar assumptions.
@@ III-A LCA group scattering
$\mathcal{F}(\xi):=\widehat{f}(\xi):=\int_{G}f(x)\overline{\xi(x)}\>\mathrm{d}\mu_{G}(x),\quad\xi\in\widehat{G}.$
In the following, we always use the unique Haar measure $\mu_{\widehat{G}}$ on $\widehat{G}$ such that the Fourier transform extends unitarily to an operator $\mathcal{F}:L^{2}(G)\to L^{2}(\widehat{G})$.
@@
$\forall h\in L^{2}(G):\leavevmode\nobreak\ \left\lVert\text{}h*\chi\text{}\right\rVert^{2}+\sum_{\psi\in\Psi}\left\lVert\text{}h*\psi\text{}\right\rVert^{2}=\left\lVert\text{}h\text{}\right\rVert^{2},$ (7)
@@
$\mathcal{H}^{(\ell)}:=L^{2}(G),\;\mathcal{L}^{(\ell)}:=\left\{\text{}\mathcal{C}_{\psi}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \psi\in\Psi\text{}\right\},\;A^{(\ell-1)}:=\mathcal{C}_{\chi},$
@@ III-B A setting-specific rate for energy decay
Let $\Gamma_{\Psi}:=\widehat{G}\setminus\bigcup_{\psi\in\Psi}\{\xi\in\widehat{G}\leavevmode\nobreak\ |\leavevmode\nobreak\ \widehat{\psi}(\xi)\neq 0\}$. We assume that $\Gamma_{\Psi}$ is a neighborhood of $\mathds{1}_{G}\in\widehat{G}$.
@@
Let $G$ be compact. If there exists $S\in\mathbb{N}$ with #$\#\operatorname{supp}(\widehat{\psi})\leq S$ ($\psi\in\Psi$), then, for all $f\in L^{2}(G)$, $N\in\mathbb{N}$,
@@
The proof of Theorem III.2 makes use of the (accessible) Fourier analysis for abelian (compact) groups, whose analogs are more involving in the nonabelian case. If $G$ is finite nonabelian, Corollary II.6 still ensures exponential energy decay, albeit with a slightly worse basis compared with the abelian case. In fact, under the low-pass condition $|\int\chi\>\mathrm{d}\mu_{G}|^{2}=1$, taking $(\lambda,\eta)=(1,\mu_{G}(G)^{-1/2}\cdot\mathds{1}_{G})$ in Corollary II.6 yields $\chi*\chi^{*}*\eta=\eta\geq\mu_{G}(G)^{-1/2}$. Since $\mu_{G}(\{e_{G}\})/\mu_{G}(G)=1/\#G$, we have, for all $f\in\mathbb{C}^{G}$, $N\in\mathbb{N}$,
@@
Thus, $\alpha(\Psi,\Gamma)<1$ if $\sup_{\psi\in\Psi}\mu_{\widehat{G}}(\Gamma\cdot\operatorname{supp}(\widehat{\psi}))<\infty$, i.e., we then have exponential energy decay globally on $L^{2}(G)$. In particular, if $G$ is discrete, exponential energy decay holds globally on $L^{2}(G)$.
@@
Theorem III.2, Remark III.3, and Theorem III.4 corroborate the experimental reports (e.g., [16, 17]) that the energy decays exponentially with increasing network depth in digital implementations of scattering CNNs, which is even reinforced by their measure-theoretic analog, Corollary II.6. However, the fact that the digital implementation of any feature extractor necessarily deals with discrete finite signals seems to play a significant role for its energy distribution. In fact, the recent paper [10] shows that the energy decay can even be arbitrarily slow in certain scattering CNNs over the Euclidean domain $\mathbb{R}^{d}$. Note that this is no contradiction to the findings in this paper, even when considering the digitial implementations of scattering CNNs as suitable discretizations of their Euclidean-domain counterparts: The basis of the exponential decay guaranteed by, e.g., Theorem III.2, converges to $1$ if $\sup_{\psi\in\Psi}\#\operatorname{supp}(\widehat{\psi})$ increases at the order of #$\#G$ as the size of the group tends to infinity.
@@
Our results contribute to the understanding of the energy distribution in deep feature extractors and, as a by-product, partially, positively answer the question of the applicability of a stability result for scattering CNNs posed in [13, Remark 7].

### v3

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. M. Getter, “Digital implementations of deep feature extractors are intrinsically informative”, 2025 International Conference on Sampling Theory and Applications (SampTA), 2025, pp. 1-6, DOI: 10.1109/SampTA64769.2025.11133552.
@@ II-A A general model for feature extraction
For every $\ell\in\mathbb{N}_{0}$, let $\mathcal{H}^{(\ell)}:=L^{2}(\mathcal{M}^{(\ell)})$ be the Lebesgue-space [^1] If there is no ambiguity, we omit the Hilbert-space index to denote the corresponding inner product $\left$$\text{}\,\cdot$$\cdot\,\text{}\right$> or norm $\left\lVert\text{}\cdot\text{}\right\rVert$. of square-integrable complex-valued functions over a measure space $\mathcal{M}^{(\ell)}:=(M^{(\ell)},\mathfrak{M}^{(\ell)},\mu^{(\ell)})$. Set $\mathcal{H}:=\mathcal{H}^{(0)}$. For a broad range of both finite- or infinite-depth neural networks, the modules for the $\ell$-th layer, $\ell\in\mathbb{N}$, of such a neural network can be described by an at most countable family of bounded linear operators $\mathcal{L}^{(\ell)}\subseteq\mathcal{B}(\mathcal{H}^{(\ell-1)},\mathcal{H}^{(\ell)})$, which forward the information from layer $\ell-1$ to layer $\ell$, a nonexpansive [^2] That is, $1$-Lipschitz, for all $f,g\in\mathcal{H}^{(\ell)}$, $\left\lVert\text{}\sigma^{(\ell)}f-\sigma^{(\ell)}g\text{}\right\rVert\leq\left\lVert\text{}f-g\text{}\right\rVert$. map $\sigma^{(\ell)}:\mathcal{H}^{(\ell)}\to\mathcal{H}^{(\ell)}$ with $\sigma^{(\ell)}(0)=0$, which allows to introduce additional complexity to the model, and a bounded linear operator $A^{(\ell)}\in\mathcal{B}(\mathcal{H}^{(\ell)})$, which allows to generate an output at depth $\ell$. Likewise, $A^{(0)}\in\mathcal{B}(\mathcal{H})$ allows to produce a first output at layer depth zero. For any input signal $f\in\mathcal{H}$, the neural network associated with these modules generates the output
$Sf:=\left\{\text{}S[p]f\penalty 10000\ :\penalty 10000\ \ell\geq 0,\penalty 10000\ p\in\mathcal{P}^{(\ell)}\text{}\right\},$
@@ II-B Energy conservation
The action of the $\ell$-th layer of a feedforward neural network [18] is typically described by $f^{(\ell)}=\sigma(W^{(\ell)}f^{(\ell-1)})$, where $W^{(\ell)}\in\mathbb{C}^{d^{(\ell)}\times d^{(\ell-1)}}$ are the (learned) weights including a potential bias term, $\sigma$ is a pointwise nonlinearity, and $f^{(\ell)}$ is the output of the $\ell$-th layer. Clearly, this fits into our framework, by setting $\mathcal{H}=\mathbb{C}^{d^{(0)}}$, $\mathcal{H}^{(\ell)}=\mathbb{C}^{d^{(\ell)}}$, $\mathcal{L}^{(\ell)}=\left\{\text{}W^{(\ell)}\text{}\right\}$, $\sigma^{(\ell)}=\sigma$, for $\ell=1,\ldots,D$, as well as $A^{(\ell)}=0_{d^{(\ell)}\times d^{(\ell)}}$ for $\ell=0,\ldots,D-1$, and $A^{(D)}=\operatorname{id}_{\mathbb{C}^{d^{(D)}}}$. Assumption II.1 then boils down to the requirement that the singular values of $W^{(\ell)}$ are $\leq 1$. Linear feedforward neural networks [19] are a special instance thereof, taking pointwise identities for $\sigma$.
@@
Further, $S:\mathcal{H}\to\bigoplus_{\ell=0}^{\infty}\bigoplus_{p\in\mathcal{P}^{(\ell)}}\mathcal{H}^{(\ell)}$ is nonexpansive with $S(0)=0$, hence norm-decreasing.
@@
The infinite sum $\sum_{\ell=N}^{\infty}\sum_{p\in\mathcal{P}^{(\ell)}}\left\lVert\text{}S[p]f\text{}\right\rVert^{2}$ provides a natural measure to quantify the information contained in the layers of depth $\geq N$. This quantity is bounded above by $W_{N}(f)$. Assuming that $Sf$ is nontrivial, truncating the (possibly infinitely deep) feature extractor after the first $N$ layers looses at most $W_{N}(f)/\left\lVert\text{}Sf\text{}\right\rVert^{2}$ percent of the total energy that is distributed across the entire network, suggesting that a fast decay of this quantity is desirable (see also [7]). Rapid energy decay across layers is also provably closely related to the stability of scattering CNNs with respect to pertubations of the input [5, Corollary 2.15], [14, Theorem 5].
@@
$=\inf_{0\neq g\in\mathcal{R}^{(N)}}\left\lVert\text{}A^{(N)}\left(\frac{g}{\left\lVert\text{}g\text{}\right\rVert}\right)\text{}\right\rVert^{2},$ (4)
@@ II-C A generic rate for energy decay
where $C^{(\ell)}:=1$ if $\mathcal{R}^{(\ell)}=\{0\}\subseteq\mathcal{H}^{(\ell)}$ and else
$C^{(\ell)}:=\inf\limits_{0\neq g\in\mathcal{R}^{(\ell)}}\penalty 10000\ \sup\limits_{(\lambda,\eta)\in E^{(\ell)}}\penalty 10000\ \left|\left<\text{}\,g/\left\lVert\text{}g\text{}\right\rVert,\eta\,\text{}\right>\right|^{2}\cdot\lambda\geq 0.$ (6)
@@
While $E^{(\ell)}=\emptyset$ may occur in the very general case, we have $E^{(\ell)}\neq\emptyset$ in many practically relevant settings. In fact, the setting is inspired by that of scattering CNNs over general measure spaces [14], where $A^{(\ell)}=H^{t}$ is the diffusion operator defined in [14, Equ. (8),(10)] for some $t>0$, so $E^{(\ell)}=\{(g(\lambda_{k})^{2t},\varphi_{k})\penalty 10000\ |\penalty 10000\ k\in I\}$ in the notation of [14].
@@
Assume that $C_{\mathcal{M}^{(\ell)}}:=\inf_{m}\penalty 10000\ \mu^{(\ell)}(m)>0$, where the infimum is taken over all sets $m\in\mathfrak{M}^{(\ell)}$ for which $\mu^{(\ell)}(m)>0$. Suppose further that $\sigma^{(\ell)}f\geq 0$ ($\mu^{(\ell)}$-a.e.) for all $f\in\mathcal{H}^{(\ell)}$ and that there exists a pair $(\lambda,\eta)\in E^{(\ell)}$ such that $\lambda^{1/2}\eta\geq C_{A^{(\ell)}}^{1/2}$ holds ($\mu^{(\ell)}$-a.e.) for a constant $C_{A^{(\ell)}}>0$. Then, we have $C^{(\ell)}=1$ if $\mathcal{R}^{(\ell)}=\{0\}\subseteq\mathcal{H}^{(\ell)}$ and else
@@
In particular, if this holds for the first $\ell=1,\ldots,N-1$ layers of the neural network, then we have, for all $f\in\mathcal{H}$,
@@
Let us now continue the discussion from above regarding the concreteness and applicability of our results. First, note that the constraint on the measure space is satisfied for the counting measure, in which case $C_{\mathcal{M}^{(\ell)}}=1$. This may even be regarded as the most practically relevant measure space, having a feature engineer in mind who works with whatever feature extractor on a digital computer. Second, there is a whole class of feature extractors, for which there exists a constant eigenfunction, $\eta\equiv C_{A^{(\ell)}}^{1/2}=\mu^{(\ell)}(M^{(\ell)})^{-1/2}$. Indeed, for scattering CNNs defined via diffusion operators as in [14, Equ. (8)], the eigenfunction $\varphi_{0}$ of $\mathcal{L}$ (in the notation of [14]) is often constant [14, Remark 3], and we have $(1,\varphi_{0})\in E^{(\ell)}$ in the context of the above corollary; see also [14, Theorems 2, 5] and [5, Section 5.1], which are based on similar assumptions.
@@ III-A LCA group scattering
$\mathcal{F}f(\xi):=\widehat{f}(\xi):=\int_{G}f(x)\overline{\xi(x)}\>\mathrm{d}\mu_{G}(x),\quad\xi\in\widehat{G}$.
In the following, we always use the unique Haar measure $\mu_{\widehat{G}}$ on $\widehat{G}$ such that the Fourier transform extends unitarily to an operator $L2(G^)\mathcal{F}$$\to L^{2}(\widehat{G})$.
@@
$\forall h\in L^{2}(G):\penalty 10000\ \left\lVert\text{}h*\chi\text{}\right\rVert^{2}+\sum_{\psi\in\Psi}\left\lVert\text{}h*\psi\text{}\right\rVert^{2}=\left\lVert\text{}h\text{}\right\rVert^{2},$ (7)
@@
$\mathcal{H}^{(\ell)}:=L^{2}(G),\;\mathcal{L}^{(\ell)}:=\left\{\text{}\mathcal{C}_{\psi}\penalty 10000\ \middle|\penalty 10000\ \psi\in\Psi\text{}\right\},\;A^{(\ell-1)}:=\mathcal{C}_{\chi},$
@@ III-B A setting-specific rate for energy decay
Let $\Gamma_{\Psi}:=\widehat{G}\setminus\bigcup_{\psi\in\Psi}\{\xi\in\widehat{G}\penalty 10000\ |\penalty 10000\ \widehat{\psi}(\xi)\neq 0\}$. We assume that $\Gamma_{\Psi}$ is a neighborhood of $\mathds{1}_{G}\in\widehat{G}$.
@@
Let $G$ be compact. If there exists $S\in\mathbb{N}$ with $\operatorname{supp}(\widehat{\psi})\leq S$ ($\psi\in\Psi$), then, for all $f\in L^{2}(G)$, $N\in\mathbb{N}$,
@@
The proof of Theorem III.2 makes use of the (accessible) Fourier analysis for abelian (compact) groups, whose analogs are more involving in the nonabelian case. If $G$ is finite nonabelian, Corollary II.6 still ensures exponential energy decay, albeit with a slightly worse basis compared with the abelian case. In fact, under the low-pass condition $|\int\chi\>\mathrm{d}\mu_{G}|^{2}=1$, taking $(\lambda,\eta)=(1,\mu_{G}(G)^{-1/2}\cdot\mathds{1}_{G})$ in Corollary II.6 yields $\eta*\chi*\chi^{*}=\eta\geq\mu_{G}(G)^{-1/2}$. Since $\mu_{G}(\{e_{G}\})/\mu_{G}(G)=1/\#G$, we have, for all $f\in\mathbb{C}^{G}$, $N\in\mathbb{N}$,
@@
[…]
@@
[…]
@@
Our results contribute to the understanding of the energy distribution in deep feature extractors and, as a by-product, partially, positively answer the question of the applicability of a stability result for scattering CNNs posed in [14, Remark 7].